OpenStudy (anonymous):

What is the sum of the first 1000 positive integers?

OpenStudy (anonymous):

See, the formula says: \[1+2+3....................+n = n(n+1)/2\] Here, n = 1000 Put in the formula and calculate the sum then...

OpenStudy (anonymous):

\[\frac{1000 \times 1001}{2} = 500 \times 1001\] Multiply this and get the answer...

OpenStudy (maheshmeghwal9):

\[S_n=\frac{n}{2}[2a+(n-1)d]\]OR\[S_n=(a+l) \times \frac{n}{2}.\]

OpenStudy (anonymous):

THere was a formula for that? I never knew that. No wonder I was getting a wrong answer. Well, now that I think about it, it might just be me being tired...It makes perfect sense now.

OpenStudy (anonymous):

500,500

OpenStudy (anonymous):

Yeah, 500*1001 = 500,500..

OpenStudy (anonymous):

FYI: It's 6:00 where I am now.

OpenStudy (anonymous):

There are three values I want to share with you: 1. Sum of First n natural Numbers: \[1+2+3+4+5...............+n = \frac{n(n+1)}{2}\] 2. Sum of squares of First n natural Numbers: \[1^2+2^2+3^2+4^2+5^2...............+n^2 = \frac{n(n+1)(2n+1)}{6}\] 3. Sum of Cubes of First n natural Numbers: \[1^3+2^3+3^3+4^3+5^3...............+n^3 = \frac{n^2(n+1)^2}{4} = [\frac{n(n+1)}{2}]^2\]

OpenStudy (anonymous):

This is the formula for a sequence which must be in Arithmetic Progression.. If value of d the common difference is same you only can then use this formula.. And in case of squares and cubes you cannot use this what so called Universal Formula @maheshmeghwal9 ..

OpenStudy (maheshmeghwal9):

oops sorry i forgt that:) but is that cube or squre GP?

OpenStudy (maheshmeghwal9):

@waterineyes

OpenStudy (anonymous):

No, it is not that.. But it is not the universal formula my dear friend..

OpenStudy (maheshmeghwal9):

k i must take back my words. thanx for capturing my fault.

OpenStudy (anonymous):

Welcome buddy..